OpenStudy (anonymous):

What is the slope intercept form of -8x+2y=4?

OpenStudy (shamil98):

-8x + 2y = 4 2y = 8x + 4 y = 4x + 2

OpenStudy (shamil98):

y = mx +b is slope intercept form btw

OpenStudy (anonymous):

Divide each term in the equation by 2. 2y2=8x2+42 Simplify the left-hand side of the equation by canceling the common factors. y=8x2+42 Simplify the equation.

OpenStudy (anonymous):

y=mx+b Using the y=mx+b formula, m=